OpenStudy (anonymous):

A 7-foot ladder is leaned against a building in such a way that the bottom of the ladder is 4 feet from the base of the wall. Find the angle formed between the ladder and the ground.

OpenStudy (nurali):

Using the cos: cos (angle formed) = (Adjacent Side) / (Hypotenuse side) Adjacent Side = Distance from ladder to wall = 4 Hypotenuse side = Length of ladder cos (angle) = 4 / 7 (angle) = cosh (4/7)=55 * cosh = cos^-1

OpenStudy (anonymous):